# StyleBooth: Image Style Editing with Multimodal Instruction

### v1

@@ 4.2 Text-Based Style Editing
Figure 6: Exemplar-based style editing comparison with 3 baselines The inputs consist of an exemplar style image for the required image style and an original image to be edited which are shown at the left-bottom corner of exemplar images. StyleBooth achieves both accurate style extraction and transformation and identical content preservation.
@@
4.2.4 Qualitative Comparisons.
Fig. 5 showcases the editing results from our method alongside those of the baseline methods. Our approach exhibits commendable performance with novel styles, such as "Picasso" and "Claymation", as well as with instructions framed in previously unseen syntax, demonstrating robust zero-shot learning abilities in text-based tasks. As previously discussed, Emu Edit falls short in adequately applying the requisite style features to the edited images, leading to a high resemblance between the original and the edited images.
Additionally, Magic Brush has a propensity to concentrate on local areas after fine-tuning on datasets specific to local editing, even when engaged in style editing tasks. In comparison to InstructPix2Pix, our method displays superior precision in style editing and mitigating artifacts. Above all, our edited images stand out as the most accurate and visually appealing among those produced by competing methods.
4.3 Exemplar-Based Style Editing
4.3.1 Baselines.
We conduct comparisons with three state-of-the-art methods, all of which are adept at extracting styles from image exemplars. RIVAL[51] is a training-free approach that harmonizes the denoising process with the exemplar inversion process through the application of interactive attention injection. By inverting the exemplar image and infusing attention features into the denoising procedure, the style of the exemplar is transferred to the resulting image. Similarly, StyleAligned[17] inverts the exemplar image into an initial noisy latent space. A pure noise sample is then denoised while sharing attention layer features with the denoising process originating from the inverted latent. To more effectively capture the style information, an adaptive normalization operation, AdaIN[21], is employed. In the VCT[8] approach, both the exemplar and source images are inverted into optimized embeddings, where style and content information are fused via classifier-free guidance[19] and attention control mechanisms. It should be noted that RIVAL and StyleAligned were originally proposed for the exemplar-based text-to-image generation task. Therefore, we supply a textual prompt derived from the original image to provide the necessary content information. To enhance content preservation, we also include results generated using ControlNet[49] with Canny-based[6] and depth-based[37] features, as utilized in their respective official implementations.
4.3.2 Comparisons.
Fig. 6 displays the results of exemplar-based style editing. RIVAL produces images that effectively capture the style features of the exemplars. Despite this, RIVAL struggles to accurately transfer the exemplar styles onto the original image, even when combined with the Canny-based ControlNet. StyleAligned demonstrates a strong ability to interpret styles. However, it inadvertently introduces unintended content features from the exemplars into the output images. When using depth-based ControlNet, StyleAligned manages to maintain the main object’s outline but at the expense of color and background details. VCT, with its specially designed and optimized content embedding, preserves a majority of the content information. Nevertheless, our method outperforms VCT in both the precision of style transfer and the overall quality of the output image. In summary, our method not only supports exemplar-based image style editing but also produces results that are both accurate and of a high aesthetic quality. Notably, our technique precisely extracts style features from the exemplars while faithfully maintaining the original image content for the style editing process.
4.4 Style Compositional Editing and Style Interpolation.
@@ 5 Conclusion
Limitations. In this work, we construct a rich dataset for style editing. However, the data construction is based on the textual descriptions of specific styles, such as watercolor, which significantly limits the number of styles. The collection of more extensive and broader editing datasets will be our future work.
Appendix 0.A Appendix
In the appendix, we present more implementation details of Data Generation in Sec. 0.A.1 including image pairs and textual instruction templates. We also provide the evaluation of our training data in Sec. 0.A.2. Finally, we showcase additional comparison results in Sec. 0.A.3.
0.A.1 Data Generation
0.A.1.1 Details of Image Pair Generation.
We use stylize prompt expansion formats provided by Fooocus[11], see the second column of Tab. 2 for examples. A placeholder of "{prompt}" is set for the original prompt. To generate $BatchA$, we use 217 various prompts to generate diverse style images. Similarly, we select 200 image captions from LAION Art[41] dataset as prompts to generate $BatchB$.  Tab. 3 shows 5 samples of prompts used for $BatchA$ and $BatchB$ generation. See the first column of Fig. 8 for additional $BatchA$ samples in different styles.
We first train a vanilla version of image editing model using the Instruct-Pix2Pix[4] training data based on a pre-trained large T2I diffusion model at 512$\times$512 resolution. Each Style Tuner and De-style Tuner is tuned on 1 NVIDIA A800-SXM4-80GB GPU for 10000 steps. We use a learning rate of 0.0001 and a small batch size of 4. The training resolution is set to 1024$\times$1024. Image pairs are randomly resized to 1×1$\times$1 × – 1.125×1.125$\times$1.125 × training resolution and then center cropped into 1×1$\times$1 ×. For usability filtering, we employ a CLIP-based metric and establish upper and lower thresholds of 0.84 and 0.2.
@@ 0.A.1 Data Generation
Table 4: Samples of text- and exemplar-based instruction templates. We utilize LLM to generate 15 templates per task. "<style>" and "<image>" are identifiers for textual style name and style exemplar image.
@@
0.A.1.2 Textual Instruction Templates.
We respectively list 5 samples from all 15 machine-generated textual instruction templates for text- and exemplar-based style editing in Tab. 4 respectively. During training, we randomly choose from these templates to keep the syntax diversity.
@@ 0.A.2 Data Evaluation
0.A.3 Additional Results
0.A.3.1 Additional Qualitative Results.
As shown in Fig. 9, we present more qualitative results in Emu Edit benchmark, including comparisons with Emu Edit[42], InstructPix2Pix[4] and MagicBrush[48]. In Fig. 10, we show additional results of exemplar-based style editing with real world images. We use two original images: the David by Michelangelo and the Eiffel Tower, five style exemplars: an animate film stage photo, a Fauvism painting by Henri Matisse, a Cubism pain-
@@ 0.A.3 Additional Results
ting by Pablo Ruiz Picasso, a post-Impressionist painting by Georges Seurat and a pixel game character. In Fig. 11, we demonstrate an editing result in a compositional style of 3 different styles. Both original image and the style exemplars are real world images. The art work "The Son of Man" by Rene Magritte is used as the original image and 3 exemplar images in different style are provided. We show the results under different scale weights.

### v2

@@ 4.2 Text-Based Style Editing
Qualitative comparisons. The editing results of StyleBooth alongside those of the baselines are shown in  Fig. 5. Our approach exhibits commendable performance with novel styles, such as "Picasso" and "Claymation", as well as with instructions framed in previously unseen syntax, demonstrating robust zero-shot learning abilities in text-based tasks. As previously discussed, Emu Edit falls short in adequately applying the requisite style features to the edited images, leading to a high resemblance between the original and the edited images. Additionally, Magic Brush has a propensity to concentrate on local areas after fine-tuning on datasets specific to local editing, even when engaged in style editing tasks. In comparison to InstructPix2Pix, our method displays superior precision in style editing and mitigating artifacts. Above all, our edited images stand out as the most accurate and visually appealing among those produced by competing methods.
4.3 Exemplar-Based Style Editing
Baselines. We conduct comparisons with three state-of-the-art methods capable of extracting styles from image exemplars. Both RIVAL[51] and StyleAligned[17] are inversion-based methods inverting exemplar images into an initial noisy latent space then infusing features via attention mechanisms. In the VCT[8] approach, both the exemplar and source images are optimized into a latent space, where style and content are fused via classifier-free guidance[18] and attention. It should be noted that RIVAL and StyleAligned are originally proposed for exemplar-based text-to-image generation task. Therefore, we give a textual prompt derived from the original image to provide the necessary content information. To further enhance their content preservation, we also include results with Canny-based[6] and depth-based[36] ControlNet[48], as utilized in their official implementations.
Comparisons. Exemplar-based style editing comparisons are displayed in  Fig. 6. Both RIVAL and StyleAligned produce images that effectively capture the style features of the exemplars. However, they struggle to keep content-consistent to original image, even when combined with the Canny-based or depth-based ControlNet. Optimizing a content embedding for each original image, VCT preserves a majority of the content information. Nevertheless, VCT falls short in terms of both style transfer precision and output image quality. In summary, our method not only supports exemplar-based image style editing but also produces results that are both accurate and of a high aesthetic quality. Notably, our technique precisely extracts style features from the exemplars while faithfully maintaining the original image content for the style editing process.
Figure 6: Exemplar-based style editing comparisons The inputs include a style exemplar and an original image displayed on the left-bottom corner of exemplar images. StyleBooth achieves both accurate style extraction and transformation and identical content preservation.
5 Style Composition and Interpolation.
@@ 6 Conclusion
Limitations. In this work, we construct a rich dataset for style editing. However, the data construction is based on the textual descriptions of specific styles, such as watercolor, which significantly limits the number of styles. Collecting more extensive editing datasets will be our future work.
Appendix A Overview
In the appendix, we present more implementation details of StyleBooth Dataset in appendix B including image pairs and textual instruction templates. We also provide the evaluation of our training data in Sec. B.4. Secondly, we explain the experiment implementation details in appendix C. We show our model’s style composition and interpolation ability and more results in Sec. 5 and  appendix D. Furthermore, we discuss social impacts and licence of assets in appendix E. Finally, we explain the details of human evaluation in appendix F.
We public the following materials:
Table 2: Shared materials and licenses in this work.
Material Link License
Homepage https://ali-vilab.github.io/stylebooth-page/ -
Code https://github.com/modelscope/scepter/tree/main/docs/en/tasks/stylebooth.md Apache-2.0
Dataset https://modelscope.cn/models/iic/stylebooth/file/view/master?fileName=datasets%252Fstylebooth_dataset.zip Apache-2.0
Model https://modelscope.cn/models/iic/stylebooth/file/view/master?fileName=models%252Fstylebooth-tb-5000-0.bin Apache-2.0
Appendix B Dataset Details.
B.1 Style Categories.
The StyleBooth dataset consists of 63 style categories and the image pair numbers of each style are shown in  Fig. 8. While some styles have fewer image pairs, the majority is relatively uniformly distributed with the number of image pairs being above 100.
Figure 8: Style distributions of StyleBooth dataset. We present the name and image pair numbers for each style. Best viewed when zoomed in.
B.2 Image Pair Construction.
We use stylize prompt expansion formats provided by Fooocus[11], see the second column of Tab. 3 for examples. A placeholder of "{prompt}" is set for the original prompt. To generate $BatchA$, we use 217 various prompts to generate diverse style images. Similarly, we select 200 image captions from LAION Art[40] dataset as prompts to generate $BatchB$. 5 samples of prompts used for $BatchA$ and $BatchB$ generation are shown in Tab. 4. See the first column of Fig. 9 for $BatchA$ samples in different styles and the fourth column for $BatchB$. To produce the derived images $BatchA^{n}$ and $BatchB^{n}$, we first train a vanilla version of image editing model using the Instruct-Pix2Pix[4] training data based on a pre-trained T2I diffusion model at 512$\times$512 resolution. Each Style Tuner and De-style Tuner is tuned on 1 NVIDIA A800-SXM4-80GB GPU for 10000 steps. We use a learning rate of 0.0001 and a small batch size of 4 and the training resolution is set to 1024$\times$1024. Image pairs are randomly resized to 1×1$\times$1 × – 1.125×1.125$\times$1.125 × training resolution and then center cropped into 1×1$\times$1 ×. For usability filtering, we employ a CLIP-based metric and establish upper and lower thresholds of 0.84 and 0.2. From Fig. 9, you can see the evolution of image quality comparing $BatchA^{2}$ to $BatchA^{1}$. The final data is the image pairs of $BatchA^{2}$ and $BatchA$ which is excellent as a style editing data.
@@ Appendix A Overview
Table 5: Samples of text- and exemplar-based instruction templates. We utilize LLM to generate 15 templates for each task. "<style>" and "<image>" are identifiers for textual style name and exemplar image respectively.
@@
B.3 Textual Instruction Templates.
We list 5 samples of machine-generated textual instruction templates for text- and exemplar-based style editing in Tab. 5 respectively. We utilize LLM to generate 15 templates per task. In these instruction templates, "<style>" and "<image>" are planted as identifiers for textual style name and exemplar image. During training, we randomly choose from these templates to keep the syntax diversity.
@@
Appendix C Experiment Implementation Details.
We utilize 8 NVIDIA A800-SXM4-80GB GPUs for our experiments. During inference, we implement classifier-free guidance for both image and text conditions. Following the recommendation in [27], we also apply a re-scaling factor of 0.5 to the outcomes. For instruction-based style editing, we fine-tune the pre-trained model[4] for 5000 steps under 0.0001 learning rate. For exemplar-based style editing, we only tune the alignment layers $W$ and the U-Net[38] decoder under the learning rate of 0.00001 for 35000 steps. Both are trained using Adam[29] optimizer. The scale weighting is only applied during inference in compositional style editing tasks. We adjust scale factors in the range of $[0.5,1.5]$.
Appendix D Additional Results.
As shown in Fig. 10, we present more qualitative results in Emu Edit benchmark, including comparisons with Emu Edit[41], InstructPix2Pix[4] and Magic Brush[47]. In Fig. 11, we show additional results of exemplar-based style editing with real world images. We use two original images: the David by Michelangelo and the Eiffel Tower, five style exemplars: an animate film stage photo, a Fauvism painting by Henri Matisse, a Cubism painting by Pablo Ruiz Picasso, a post-Impressionist painting by Georges Seurat and a pixel game character. In Fig. 12, we demonstrate an editing result in a compositional style of 3 different styles. Both original image and the style exemplars are real world images. The art work "The Son of Man" by Rene Magritte is used as the original image and 3 exemplar images in different style are provided. We show the results under different scale weights.
@@
Appendix E Discussions.
E.1 Social Impacts.
StyleBooth enables the facile editing of images through simple instructions, providing users with multimodal input options to manipulate images according to their preferences. This approach leads to some positive social impacts, as it allows users to achieve style transfer effortlessly without the need for professional editing tools, substantially enhancing productivity. Moreover, users can explore more liberating style options using text prompts or reference images, thereby offering a creatively enriched editing tool. However, given the inherent risks associated with generative image synthesis, such as malicious use and dissemination, it is imperative to incorporate additional safeguards during the development of such systematic tools.
E.2 License of Assets.
For baselines, Instruction-Pix2Pix[4] inherits this license as it is built upon Stable Diffusion. Magic Brush[47] are released under Creative Commons Attribution 4.0 License. RIVAL[51], StyleAligned[17] and VCT[8] are under Apache-2.0 license.
For datasets, Emu Edit[41] and LAION Art[40] are under Creative Commons Attribution 4.0 License. According to Stable Diffusion-XL[1], which is under Open RAIL++-M License, we have the right to distribute the generated images for research purpose.
Appendix F Human Evaluation.
We conducted human evaluation to assess the baseline comparisons. We distributed anonymous questionnaires with 5 questions for each editing case, along with instruction, original image, edit result from Method 1 and Method 2 as showed in Fig. 13. The first 2 questions are about whether styles of the 2 editing results are correct and match instructions. The third and fourth questions estimate the content consistency of the 2 methods. At last, the candidates were asked to choice the better one between these 2 methods.
Figure 13: Screenshots of our anonymous questionnaires for human evaluation.